OpenStudy (anonymous):

In ∆ABC, B is a right angle and tan A = 13/16. What is mC ?

OpenStudy (anonymous):

87° 39° 51° 49°

OpenStudy (anonymous):

those are my possible answers

OpenStudy (anonymous):

well do you know the phrase SOH CAH TOA???

OpenStudy (anonymous):

well tan = opposite / adjacent

OpenStudy (anonymous):

so you will get the number then you will use \[\tan^{-1} (13/16) = your degree\]

OpenStudy (anonymous):

the ans is 51°

OpenStudy (anonymous):

nope @FriedRice its mC = tan^−1(16/13)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

@CO_oLBoY why do you think its that??

OpenStudy (anonymous):

coz tanA = 13/16 and B is a right angle and if u make a triangle then u'll know that tan C = 16/13

OpenStudy (anonymous):

well c is the angle so it should be opposite/adjacent

OpenStudy (anonymous):

not adjacent / opposite